March 11, 2015 Calibration
GlueX Calibration Meeting
Wednesday, March 11, 2015
11:00 am, EDT
JLab: CEBAF Center, F326
Connection Using Bluejeans
- To join via Polycom room system go to the IP Address: 220.127.116.11 (bjn.vc) and enter the meeting ID: 630804895.
- To join via a Web Browser, go to the page  https://bluejeans.com/630804895.
- To join via phone, use one of the following numbers and the Conference ID: 630804895
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4
- More information on connecting to bluejeans is available.
- Action Items / Subdetectors
- Update FCAL gains using LED results, and electrons if possible (Matt S./Adesh)
- Update BCAL gains using pi0 calibrations (Will M.)
- Extract CDC gas mix with plugin (Naomi)
- Update FDC alignment, if possible (Lubomir)
- Add TOF calibrations performed with cosmics (Sasha O.)
- Update timewalk and propogation length corrections for Start Counter (Eric)
- Update timing constants (Mike S.)
- Make plans to update the offline monitoring (Sean)
Attending: Sean (NU); Simon, Kei, Adesh, Eric, Paul, Alex, Justin, Mahmoud, Mark I., Elton, Eugene, Mike S. (JLab); Curtis, Naomi (CMU); Zisis (Regina)
- Paul announced that we can make small EVIO skims during monitoring runs to help with calibrations. These should be small, targeted skims, of a few percent or less of the data. Currently, a BCAL pi0 skim requested by Will McGinley is planned to be run.
- Mark I. tagged sim-recon in preparation for a new release. He changed the version numbering system so that this release is numbered 1.0.0. Please aim any complaints or flames at him.
- FCAL: Adesh reported the first pass of improved gains using electron has been performed. These affect ~600 blocks near the beampipe. He has passed these on to John Zarling to check what improvement in the pi0 peak is seen, and if they would aid in doing a pi0 calibration of the inner portion of this detector. He and Sean will discuss incorporating these into the CCDB by this Friday.
- BCAL: Will McGinley reported that updated BCAL gains were added to the CCDB yesterday. They reduce the pi0 resolution from sigma = 19 MeV -> 12 MeV. He is ready to study and update these with data from the upcoming run. Sean mentioned that it would be good to see a table of pi0 mass and resoltuion as a function of minimum cluster energy, which could help guide other calibration efforts.
- CDC: Naomi has been working on a plugin to monitor the gas mixture. This can be done using the maximum fADC amplitudes or the drift time spectrum. She has a program working to fit the drift time spectrum from one straw and running over data, and is looking to monitor ~10-20 straws together, to speed up the procedure. Selecting a good small set of straws is a time-consuming task.
- FDC: The FDC is ready to run, aside from a bad crate which is being studied, and an early goal of the spring run is to collect more data with the magnetic field off to be used for aligment. Sean will check with Lubomir about if his current alignment constants are ready to be added.
- TOF: Several items were discussed:
- In previous TOF group meetings, Sasha O. described some technical difficulties he was having with his global calibration fit procedure, and agreed to add constants based on cosmic ray data to the CCDB. However, some changes to the hit factory were needed as well, and he decided to wait until the commissioning branch was merged with the trunk. Sean will check with him to make sure that we are on track to have these constants in by Friday.
- Others have been looking at various aspects of these calibrations - Beni, Simon, Sean - and have codes more-or-less ready. They will discuss and produce a plan on how to proceed on Thursday.
- Paul mentioned that the tracks match TOF hits only ~60-65% of the time due to the lack of updated calibrations. Elton suggested loosening some of the matching parameters, and Paul agreed to look into the effects of doing that.
- Start Counter: Eric reported that a new set of timewalk corrections is ready, and that he recently finished testing a propagation time correction plugin. These constants should be ready by Friday. The framework to use these constants has been put in the PID library.
- Mike S. reported that he has run into problems running his procedure with large data sets on batch farm. The problem is essentially that after running over the data, there is often one thread that does not properly stop. He has worked extensively with David L. and debugging tools to get to the bottom of this problem with no luck. He is currently working on restructuring how he processes the data. His current constants from run 2209 should be good for the FCAL+BCAL data taken in the previous run and are a good start for this run, and agreed to add them to the CCDB by Friday.
- Trigger / Reference Time: Nathan reported that he is looking at calibrating the fine PS detector again the trigger. Improvement in trigger timing requires a reference time so that the TDC data can be analyzed.
- Beni said that the hardware for the reference times has been installed but generally needs to be checked. He did check it out for the FDC crate, and it looks very good
- Paul and Beni will discuss how to propagate this info into the analysis framework and how to best monitor it.
- Scheduled run of offline monitoring is by close-of-business on Friday, pending TOF calibration decision.
- Mark made a general remark that at this point we might want to focus on calibrating the detector with an eye towards future data, rather than doing a detailed calibration of all the different conditions of the data taken in the fall. There was general assent. Sean also mentioned that plugins should be checked into SVN, and diagnostic output of the calibration plugins could also be worked into the offline monitoring.